# Active Diffusion Subsampling

### v1

@@ 4.1 Active Diffusion Subsampling
1 $t=0;\quad A_{0}=A_{\texttt{init}};\quad y_{0}={\mathbf{U}}(A_{0})f({\mathbf{x}})+{\mathbf{n}}_{0};\quad\{{\mathbf{x}}^{(i)}_{T}\sim\mathcal{N}(\textbf{0},{\mathbf{I}})\}_{i=0}^{N_{p}-1}$
2 for $\tau=T$ to $1$ do
@@ 5            $\hat{\textbf{s}}\leftarrow\textbf{s}_{\theta}({\mathbf{x}}^{(i)}_{\tau},\tau)$
6            $\hat{{\mathbf{x}}}^{(i)}_{0}\leftarrow\mathcal{D}_{\tau}({\mathbf{x}}^{(i)}_{\tau})=\frac{1}{\sqrt{\bar{\alpha}_{\tau}}}({\mathbf{x}}^{(i)}_{\tau}+(1-\bar{\alpha}_{\tau})\hat{\textbf{s}})$
            $\hat{{\mathbf{y}}}^{(i)}\leftarrow f(\hat{{\mathbf{x}}}^{(i)}_{0})$ // Estimate the full measurement
7
@@ 8            ${\mathbf{z}}\sim\mathcal{N}(\textbf{0},{\mathbf{I}})$
9            ${\mathbf{x}}^{(i)^{\prime}}_{\tau-1}\leftarrow\frac{\sqrt{\alpha_{\tau}}(1-\bar{\alpha}_{\tau-1})}{1-\bar{\alpha}_{\tau}}{\mathbf{x}}^{(i)}_{\tau}+\frac{\sqrt{\bar{\alpha}_{\tau-1}}\beta_{\tau}}{1-\bar{\alpha}_{\tau}}\hat{{\mathbf{x}}}^{(i)}_{0}+\tilde{\sigma}_{\tau}{\mathbf{z}}$
10            ${\mathbf{x}}^{(i)}_{\tau-1}\leftarrow{\mathbf{x}}^{(i)^{\prime}}_{\tau-1}-\zeta\nabla_{{\mathbf{x}}^{(i)}_{\tau}}||{\mathbf{y}}_{t}-{\mathbf{U}}(A_{t})\hat{{\mathbf{y}}}^{(i)}||^{2}_{2}$
@@ 14             $t\leftarrow t+1$
15            $a^{*}_{t}=\operatorname*{arg\,max}_{a_{t}}\left[\sum_{i}^{N_{p}}\log\sum_{j}^{N_{p}}\exp\left\{\frac{\sum\limits_{l\in a_{t}}(\hat{{\mathbf{y}}}_{l}^{(i)}-\hat{{\mathbf{y}}}_{l}^{(j)})^{2}}{2\sigma_{\mathbf{y}}^{2}}\right\}\right]$
@@ 16            $A_{t}=A_{t-1}\cup a^{*}_{t}$
            ${\mathbf{y}}_{t}={\mathbf{U}}(A_{t})f({\mathbf{x}})+{\mathbf{n}}_{t}$ // Acquire new measurements
17
@@
return: $\hat{{\mathbf{x}}}_{0}=\frac{1}{N_{p}}\sum_{i}^{N_{p}}\hat{{\mathbf{x}}}^{(i)}_{0}$ // Return the posterior mean
22
@@
ADS operates by running a reverse diffusion process for a batch $\{{\mathbf{x}}^{(i)}_{\tau}\},i\in{0,...,N_{p}}$ guided by an evolving set of measurements $\{{\mathbf{y}}_{t}\}_{t=0}^{\mathcal{T}}$ which are revealed to it through subsampling actions taken at reverse diffusion steps satisfying $\tau\in S$, where $S$ is a subsampling schedule. We refer to the elements of this batch $\{{\mathbf{x}}^{(i)}_{\tau}\}$ as particles in the data space, as they implicitly track a belief distribution over the true, fully-denoised ${\mathbf{x}}={\mathbf{x}}_{0}$ throughout reverse diffusion. These particles are used to compute estimates of uncertainty about ${\mathbf{x}}$, which ADS aims to minimize by choosing actions $a_{t}$ that maximize the mutual information between ${\mathbf{x}}$ and ${\mathbf{y}}_{t}$ given $a_{t}$, as explained in Section 3.1. The remainder of the section describes how this is achieved through (i) employing running estimates of ${\mathbf{x}}_{0}$ given by $\mathcal{D}_{\tau}({\mathbf{x}}_{\tau})$, (ii) modelling assumptions on the measurement entropy, and (iii) computational advantages afforded by the subsampling operator.
@@
ADS follows an information-maximizing policy, selecting measurements $a_{t}^{*}=\operatorname*{arg\,max}_{a_{t}}[I({\mathbf{y}};{\mathbf{x}}\mid A_{t},{\mathbf{y}}_{t-1})]$. Assuming a measurement model with additive noise ${\mathbf{y}}_{t}={\mathbf{U}}(A_{t})f({\mathbf{x}})+{\mathbf{n}}_{t}$, the posterior entropy term $H({\mathbf{y}}_{t}\mid{\mathbf{x}},A_{t},{\mathbf{y}}_{t-1})$ in the mutual information is unaffected by the choice of action $a_{t}$: it is solely determined by the noise component ${\mathbf{n}}_{t}$. This simplifies the policy function, leaving only the marginal entropy term $H({\mathbf{y}}_{t}\mid A_{t},{\mathbf{y}}_{t-1})$ to maximize. This entropy term (Equation 2) can be computed as an expectation of $\log p({\mathbf{y}}_{t}\mid A_{t},{\mathbf{y}}_{t-1})$, or the expected distribution of measurements ${\mathbf{y}}_{t}$ given the observations so far, and a subsampling matrix $A_{t}$. We model these probability distributions in Equation 8 as mixtures of $N_{p}$ isotropic Gaussians, with means set as estimates of future measurements under possible actions, and variance $\sigma^{2}_{y}{\mathbf{I}}$. The means $\hat{{\mathbf{y}}}^{(i)}_{t}={\mathbf{U}}(A_{t})f(\hat{{\mathbf{x}}}^{(i)}_{0})$ are computed by applying the forward model to posterior samples $\hat{{\mathbf{x}}}^{(i)}_{0}\sim p({\mathbf{x}}\mid{\mathbf{y}}_{t-1})$, which are estimated using the batch of partially denoised particles ${\mathbf{x}}^{(i)}_{\tau}$ via $\hat{{\mathbf{x}}}^{(i)}_{0}=\mathcal{D}_{\tau}({\mathbf{x}}^{(i)}_{\tau})$, yielding the GMM:
$p({\mathbf{y}}_{t}\mid A_{t},{\mathbf{y}}_{t-1})=\sum_{i=0}^{N}{\mathbf{w}}_{i}\mathcal{N}(\hat{{\mathbf{y}}}^{(i)}_{t},\sigma^{2}_{\mathbf{y}}{\mathbf{I}})$ (8)
@@
Under this isotropic Gaussian Mixture Model, the marginal entropy can be approximated as follows, as given by Hershey and Olsen (2007):
@@
$H({\mathbf{y}}_{t}\mid A_{t},\hat{{\mathbf{y}}}_{0:t-1})\approx\text{constant}+\sum_{i}^{N_{p}}w_{i}\log\sum_{j}^{N_{p}}w_{j}\exp\left\{\frac{||\hat{{\mathbf{y}}}_{t}^{(i)}-\hat{{\mathbf{y}}}_{t}^{(j)}||_{2}^{2}}{2\sigma_{y}^{2}}\right\}$ (9)
@@
Figure 3: Illustration of a single action selection based on a set of partially-denoised particles $\{{\mathbf{x}}_{\tau}^{(i)}\}$.
@@
This can be further simplified for element-wise subsampling, in which $a_{t}$ represents just a single index, and the squared L2 norm simplifies to a squared error. In this case, denoting as $\mathcal{E}_{i,j}$ the element-wise Gaussian error matrix between particles $\hat{{\mathbf{y}}}^{(i)}$ and $\hat{{\mathbf{y}}}^{(j)}$, we can compute the policy as follows:
$\mathcal{E}_{i,j}=\exp\left\{\frac{(\hat{{\mathbf{y}}}^{(i)}-\hat{{\mathbf{y}}}^{(j)})^{2}}{2\sigma_{y}^{2}}\right\};\quad\mathcal{E}=\sum_{i}^{N_{p}}\log\sum_{j}^{N_{p}}\mathcal{E}_{i,j};\quad a^{*}_{t}=\operatorname*{arg\,max}_{a_{t}}\mathcal{E}[a_{t}]$ (11)
Where the $[a_{t}]$ operator selects elements indexed by $a_{t}$ from matrix $\mathcal{E}$.
4.2 Extension of ADS to Latent Diffusion Models
Current state-of-the-art diffusion models are gravitating towards a latent design, which has proven to work effectively on high resolution image synthesis (Rombach et al., 2022) and inversion (Rout et al., 2024). Latent diffusion models (LDMs) involve a two-stage training process. First an autoencoder is trained such that $\hat{{\mathbf{x}}}=D({\mathbf{z}})=D(E({\mathbf{x}}))$, with encoder $E$, decoder $D$ and ${\mathbf{z}}$ denoting the latent variable. Projecting the data into a latent space allows more effective and computationally efficient training in the semantically compressed space. Similarly, by leveraging LDMs, ADS can be scaled to unlock new domains requiring higher resolution. Recently, guidance methods similar to (7) have been extended for LDMs (Rout et al., 2024, 2023). We here show that ADS too can be straightforwardly extended with LDMs using existing guidance methods: estimates of fully sampled measurement particles $\hat{{\mathbf{y}}}^{(i)}$ are now computed by first decoding latent variables to the data-space i.e. $p(\hat{{\mathbf{y}}}^{(i)}|\hat{{\mathbf{x}}}_{0}=D(\mathbb{[}{\mathbf{z}}_{0}|{\mathbf{z}}_{\tau}]))$. We include several examples using a pretrained Stable Diffusion (Rombach et al., 2022) model out-of-the-box in Appendix A.5 to demonstrate this.
@@ 5 Experiments
5.1 Evaluation of Subsampling Strategies
@@
In order to evaluate the effectiveness of the Maximum Entropy subsampling strategy employed by ADS, we compare it to two baseline subsampling strategies on the task of reconstructing images of digits from the MNIST dataset LeCun et al. (1998). To this end, a Denoising Diffusion Implicit Model (DDIM) Song et al. (2020a) was trained on the MNIST training dataset, resized to $32\times 32$ pixels. See Appendix A.2.1 for further details on training and architecture. Using this trained DDIM, each subsampling strategy was used to reconstruct 500 unseen samples from the MNIST test set for various subsampling rates. Both pixel-based and line-based subsampling were evaluated, where line-based subsampling selects single-pixel-wide columns. The measurement model is thus ${\mathbf{y}}_{t}={\mathbf{U}}(A_{t}){\mathbf{x}}$, as there is no measurement noise or measurement transformation, i.e. $f({\mathbf{x}})={\mathbf{x}}$. The baseline subsampling strategies used for comparison were (i) Random subsampling selects measurement locations from a uniform categorical distribution without replacement, and (ii) Data Variance subsampling selects measurement locations without replacement from a categorical distribution in which the probability of a given location is proportional to the variance across that location in the training set, leading to a data-driven but fixed design strategy. Inference was performed using Diffusion Posterior Sampling for measurement guidance, with guidance weight $\zeta=1$ and $T=1000$ reverse diffusion steps. For ADS, measurements were taken at regular intervals in the window $[0,800]$, with $16$ particles. For the fixed sampling strategies, the subsampling masks were set a priori, such that all diffusion steps are guided by the measurements, as is typical in inverse problem solving with diffusion models. The results to this comparison are given in Table 1, and illustrated by Figure 4. It is clear from these results that ADS outperforms fixed mask baselines, most notably in comparison with data-variance sampling: for pixel-based sampling, we find that maximum entropy sampling with a budget of $100$ pixels outperforms data variance sampling with a budget of $250$ pixels, i.e. actively sampling the measurements is as good as having 2.5×2.5$\times$2.5 × the number of measurements with the data-variance strategy. We also find that the standard deviation of the reconstruction errors over the test set is significantly lower for $25\%$ and $50\%$ subsampling rates (typically $\sim 2$-3×3$\times$3 × than baselines, leading to more reliable reconstructions). In aggregate, these results motivate the use of ADS in applications in which naïve or fixed subsampling strategies are being used for inverse problem solving with diffusion models.
@@
No. Pixels Sampled (of total) Random Data Variance ADS (Ours)
10 (0.97%) 0.231 (0.052) 0.197 (0.048) 0.207 (0.054)
25 (2.44%) 0.190 (0.048) 0.125 (0.037) 0.124 (0.055)
50 (4.88%) 0.140 (0.041) 0.067 (0.026) 0.042 (0.030)
100 (9.76%) 0.074 (0.029) 0.034 (0.014) 0.011 (0.005)
250 (24.41%) 0.024 (0.009) 0.015 (0.006) 0.007 (0.002)
500 (48.82%) 0.011 (0.003) 0.008 (0.003) 0.007 (0.002)
No. Lines Sampled (of total) Random Data Variance ADS (Ours)
2 (6.25%) 0.197 (0.059) 0.152 (0.045) 0.148 (0.042)
4 (12.5%) 0.143 (0.054) 0.086 (0.034) 0.071 (0.033)
8 (25%) 0.073 (0.033) 0.037 (0.021) 0.019 (0.009)
16 (50%) 0.022 (0.014) 0.013 (0.007) 0.008 (0.003)
24 (75%) 0.010 (0.004) 0.0076 (0.0025) 0.0074 (0.0023)
Table 1: Mean and (standard deviation) for the Mean Absolute Error ($\downarrow$) in reconstruction of MNIST samples, using pixel- and line-based subsampling.
5.2 Application in MRI Acceleration

### v2

@@ 4.1 Active Diffusion Subsampling
1 $t=0;\quad A_{0}=A_{\texttt{init}};\quad{\mathbf{y}}_{0}={\mathbf{U}}(A_{0})f({\mathbf{x}})+{\mathbf{n}}_{0};\quad\{{\mathbf{x}}^{(i)}_{T}\sim\mathcal{N}(\textbf{0},{\mathbf{I}})\}_{i=0}^{N_{p}-1}$
2 for $\tau=T$ to 00 do
       // Batch process in parallel for efficient inference
@@ 4             $\hat{\textbf{s}}\leftarrow\textbf{s}_{\theta}({\mathbf{x}}^{(i)}_{\tau},\tau)$
            $\hat{{\mathbf{x}}}^{(i)}_{0}\leftarrow\mathcal{D}_{\tau}({\mathbf{x}}^{(i)}_{\tau})=\frac{1}{\sqrt{\bar{\alpha}_{\tau}}}({\mathbf{x}}^{(i)}_{\tau}+(1-\bar{\alpha}_{\tau})\hat{\textbf{s}})$ // Calculate Tweedie estimate
5
            $\hat{{\mathbf{y}}}^{(i)}\leftarrow f(\hat{{\mathbf{x}}}^{(i)}_{0})$ // Simulate full measurement
6
@@ 7            ${\mathbf{z}}\sim\mathcal{N}(\textbf{0},{\mathbf{I}})$
            ${\mathbf{x}}^{(i)^{\prime}}_{\tau-1}\leftarrow\frac{\sqrt{\alpha_{\tau}}(1-\bar{\alpha}_{\tau-1})}{1-\bar{\alpha}_{\tau}}{\mathbf{x}}^{(i)}_{\tau}+\frac{\sqrt{\bar{\alpha}_{\tau-1}}\beta_{\tau}}{1-\bar{\alpha}_{\tau}}\hat{{\mathbf{x}}}^{(i)}_{0}+\tilde{\sigma}_{\tau}{\mathbf{z}}$ // Reverse diffusion step
8
            ${\mathbf{x}}^{(i)}_{\tau-1}\leftarrow{\mathbf{x}}^{(i)^{\prime}}_{\tau-1}-\zeta\nabla_{{\mathbf{x}}^{(i)}_{\tau}}||{\mathbf{y}}_{t}-{\mathbf{U}}(A_{t})\hat{{\mathbf{y}}}^{(i)}||^{2}_{2}$ // Data fidelity step
9
10
@@ 13             $t\leftarrow t+1$
            // Find the group of pixels with maximum entropy across simulated measurements
14             $a^{*}_{t}=\operatorname*{arg\,max}_{a_{t}}\left[\sum\limits_{i}^{N_{p}}\log\sum\limits_{j}^{N_{p}}\exp\left\{\frac{\sum\limits_{l\in a_{t}}(\hat{{\mathbf{y}}}_{l}^{(i)}-\hat{{\mathbf{y}}}_{l}^{(j)})^{2}}{2\sigma_{\mathbf{y}}^{2}}\right\}\right]$
@@ 15            $A_{t}=A_{t-1}\cup a^{*}_{t}$
            ${\mathbf{y}}_{t}={\mathbf{U}}(A_{t})f({\mathbf{x}})+{\mathbf{n}}_{t}$ // Acquire new measurements of the true ${\mathbf{x}}$
16
@@
ADS (Algorithm 1) operates by running a reverse diffusion process generating a batch $\{{\mathbf{x}}^{(i)}_{\tau}\},i\in{0,...,N_{p}}$ of possible reconstructions of the target signal guided by an evolving set of measurements $\{{\mathbf{y}}_{t}\}_{t=0}^{\mathcal{T}}$ which are revealed to it through subsampling actions taken at reverse diffusion steps satisfying $\tau\in S$, where $S$ is a subsampling schedule, or set of diffusion time steps at which to acquire new measurements. For example, in an application in which one would like to acquire $10$ measurements, $\{{\mathbf{y}}_{t}\}_{t=0}^{9}$, across a diffusion process consisting of $T=100$ steps, one might choose a linear subsampling schedule $S=\{0,10,20,...,90\}$. Then, during the diffusion process, if the diffusion step $\tau$ is an element of $S$, a new measurement will be acquired, revealing it to the diffusion model. These measurements could be, for example, individual pixels or groups of pixels. We refer to the elements of the batch of reconstructions $\{{\mathbf{x}}^{(i)}_{\tau}\}$ as particles in the data space, as they implicitly track a belief distribution over the true, fully-denoised ${\mathbf{x}}={\mathbf{x}}_{0}$ throughout reverse diffusion. These particles are used to compute estimates of uncertainty about ${\mathbf{x}}$, which ADS aims to minimize by choosing actions $a_{t}$ that maximize the mutual information between ${\mathbf{x}}$ and ${\mathbf{y}}_{t}$ given $a_{t}$. Figure 3 illustrates this action selection procedure. The remainder of the section describes how this is achieved through (i) employing running estimates of ${\mathbf{x}}_{0}$ given by $\mathcal{D}_{\tau}({\mathbf{x}}_{\tau})$, (ii) modeling assumptions on the measurement entropy, and (iii) computational advantages afforded by the subsampling operator.
@@
ADS follows an information-maximizing policy, selecting measurements $a_{t}^{*}=\operatorname*{arg\,max}_{a_{t}}[I({\mathbf{y}};{\mathbf{x}}\mid A_{t},{\mathbf{y}}_{t-1})]$. Assuming a measurement model with additive noise ${\mathbf{y}}_{t}={\mathbf{U}}(A_{t})f({\mathbf{x}})+{\mathbf{n}}_{t}$, the posterior entropy term $H({\mathbf{y}}_{t}\mid{\mathbf{x}},A_{t},{\mathbf{y}}_{t-1})$ in the mutual information is unaffected by the choice of action $a_{t}$: it is solely determined by the noise component ${\mathbf{n}}_{t}$. This simplifies the policy function, leaving only the marginal entropy term $H({\mathbf{y}}_{t}\mid A_{t},{\mathbf{y}}_{t-1})$ to maximize. This entropy term (Equation 2) can be computed as the expectation of $\log p({\mathbf{y}}_{t}\mid A_{t},{\mathbf{y}}_{t-1})$, or the expected distribution of measurements ${\mathbf{y}}_{t}$ given the observations so far, and a subsampling matrix $A_{t}$. We approximate the true measurement posterior in Equation 8 as a mixture of $N_{p}$ isotropic Gaussians, with means set as estimates of future measurements under possible actions, and variance $\sigma^{2}_{y}{\mathbf{I}}$. The means $\hat{{\mathbf{y}}}^{(i)}_{t}={\mathbf{U}}(A_{t})f(\hat{{\mathbf{x}}}^{(i)}_{0})$ are computed by applying the forward model to posterior samples $\hat{{\mathbf{x}}}^{(i)}_{0}\sim p({\mathbf{x}}\mid{\mathbf{y}}_{t-1})$, which are estimated using the batch of partially denoised particles ${\mathbf{x}}^{(i)}_{\tau}$ via $\hat{{\mathbf{x}}}^{(i)}_{0}=\mathcal{D}_{\tau}({\mathbf{x}}^{(i)}_{\tau})$, yielding:
$p({\mathbf{y}}_{t}\mid A_{t},{\mathbf{y}}_{t-1})=\sum_{i=0}^{N_{p}}w_{i}\mathcal{N}(\hat{{\mathbf{y}}}^{(i)}_{t},\sigma^{2}_{\mathbf{y}}{\mathbf{I}})$ (8)
@@
Note the parameter $\sigma_{y}^{2}$, which controls the width of the Gaussians . The entropy of this isotropic Gaussian Mixture Model can then be approximated as follows, as given by Hershey & Olsen (2007):
@@
$H({\mathbf{y}}_{t}\mid A_{t},{\mathbf{y}}_{t-1})\approx\text{constant}+\sum_{i}^{N_{p}}w_{i}\log\sum_{j}^{N_{p}}w_{j}\exp\left\{\frac{||\hat{{\mathbf{y}}}_{t}^{(i)}-\hat{{\mathbf{y}}}_{t}^{(j)}||_{2}^{2}}{2\sigma_{\mathbf{y}}^{2}}\right\}$ (9)
Intuitively, this approximation computes the entropy of the mixture of Gaussians by summing the Gaussian error between each pair of means, which grows with distance according to the hyperparameter $\sigma_{\mathbf{y}}^{2}$. This hyperparameter $\sigma_{\mathbf{y}}^{2}$ can therefore be tuned to control the sensitivity of the entropy to outliers.
@@
[…]
@@ 5 Experiments
We evaluate our method with three sets of experiments, covering a variety of data distributions and application domains. The first two experiments, in Sections 5.1 and 5.2, evaluate the proposed Maximum-Entropy subsampling strategy employed by ADS against baseline strategies, keeping the generative model fixed to avoid confounding. Next, in Section 5.3, we evaluate the model end-to-end on both sampling and reconstruction by applying it in the real-world task of MRI acceleration with the fastMRI dataset, and comparing it to existing supervised approaches.
5.1 MNIST
@@
In order to evaluate the effectiveness of the Maximum Entropy subsampling strategy employed by ADS, we compare it to two baseline subsampling strategies on the task of reconstructing images of digits from the MNIST dataset (LeCun et al., 1998). To this end, a diffusion model was trained on the MNIST training dataset, resized to $32\times 32$ pixels. See Appendix A.2.1 for further details on training and architecture. Using this trained diffusion model, each subsampling strategy was used to reconstruct 500 unseen samples from the MNIST test set for various subsampling rates. Both pixel-based and line-based subsampling were evaluated, where line-based subsampling selects single-pixel-wide columns. The measurement model is thus ${\mathbf{y}}_{t}={\mathbf{U}}(A_{t}){\mathbf{x}}$, as there is no measurement noise or measurement transformation, i.e. $f({\mathbf{x}})={\mathbf{x}}$. The baseline subsampling strategies used for comparison were as follows: (i) Random subsampling selects measurement locations from a uniform categorical distribution without replacement, and (ii) Data Variance subsampling selects measurement locations without replacement from a categorical distribution in which the probability of a given location is proportional to the variance across that location in the training set. This can therefore be seen as a data-driven but fixed design strategy.
Inference was performed using Diffusion Posterior Sampling for measurement guidance, with guidance weight $\zeta=1$ and $T=1000$ reverse diffusion steps. For ADS, measurements were taken at regular intervals in the window $[0,800]$, with $N_{p}=16$ particles and $\sigma_{{\mathbf{y}}}=10$. For the fixed sampling strategies, the subsampling masks were set a priori, such that all diffusion steps are guided by the measurements, as is typical in inverse problem solving with diffusion models. The results to this comparison are illustrated by Figure 4, with numbers provided in Appendix A.3. We use Mean Absolute Error as an evaluation metric since MNIST consists of single channel brightness values. It is clear from these results that ADS outperforms fixed mask baselines, most notably in comparison with data-variance sampling: for pixel-based sampling, we find that maximum entropy sampling with a budget of $100$ pixels outperforms data variance sampling with a budget of $250$ pixels, i.e. actively sampling the measurements is as good as having 2.5×2.5$\times$2.5 × the number of measurements with the data-variance strategy. We also find that the standard deviation of the reconstruction errors over the test set is significantly lower for $25\%$ and $50\%$ subsampling rates (typically $\sim 2$-3×3$\times$3 × than baselines, leading to more reliable reconstructions). Further experiments on MNIST are carried out in Appendix A.4, where ADS is compared to Active Deep Probabilistic Subsampling Van Gorp et al. (2021), an end-to-end supervised method for active subsampling.
@@
5.2 CelebA
In order to evaluate ADS on a natural image dataset, a diffusion model was trained on the CelebA (Liu et al., 2015) training dataset at 128 $\times$ 128 resolution (see Appendix A.2.2 for training details). ADS was then benchmarked on N=100 samples from the CelebA test set against DPS with baseline sampling strategies. This evaluation was carried out for a number sampling rates $|S|\in\{50,100,200,300,500\}$. The measurement scheme employed here samples ‘boxes’ of size 4$\times$4 pixels from the image. The number of diffusion steps taken during inference was chosen based on sampling rate, with 400 steps for $|S|<300$, 600 steps for $|S|=500$. In each case, for sampling rate $|S|$ the sampling window $[10,|S|+10]$ was used. These sampling windows were chosen empirically, finding that some initial unguided reverse diffusion steps help create a better initial estimate of the posterior. $N_{p}=16$ particles with $\sigma_{{\mathbf{y}}}=1$ were used to model the posterior distribution. The peak signal-to-noise ratio (PSNR) values between the posterior mean and test samples are provided in Figure 5(a) (a) and Table 1, with further metrics, including metrics across individual posterior samples, are provided in Appendix A.5. It is clear in these results that ADS outperforms baseline sampling strategies with DPS across a number of sampling rates. Note, for example, that ADS with 200 measurements achieves a higher PSNR than random-mask DPS with 300 measurements. In Figure 5(b) a selection of samples are provided to exemplify the mask designs produced by ADS. Note here that the masks focus on important features such as facial features in order to minimize the posterior entropy, leading to higher information gain about the target, and ultimately better recovery.
Figure 5: In (a), PSNR ($\uparrow$) scores for ADS vs DPS with random and data variance sampling on N=100 unseen samples from the CelebA dataset are plotted, for increasing numbers of measurements. A measurement here is a $4\times 4$ box of pixels. (b) shows some examples of ADS inference versus DPS with random measurements on the CelebA dataset from the evaluation with 200 boxes sampled.
# Samples (Max %) Random DPS Data Variance DPS ADS
50 (4.88%) 18.099 (0.019) 17.747 (0.018) 18.932 (0.024)
100 (9.76%) 20.580 (0.019) 19.892 (0.017) 22.725 (0.024)
200 (19.53%) 23.555 (0.020) 22.895 (0.018) 27.954 (0.026)
300 (29.29%) 25.919 (0.019) 24.779 (0.019) 31.483 (0.027)
500 (48.82%) 29.123 (0.018) 27.512 (0.018) 35.630 (0.027)
Table 1: PSNR ($\uparrow$) on test samples from the CelebA dataset. Next to each # Samples $|S|$ is a ’Max %’ indicating the maximum % of a $128\times 128$ image that $|S|$ $4\times 4$ boxes could cover.
5.3 MRI Acceleration